# REVEAL-IT: REinforcement learning with Visibility of Evolving Agent poLicy for InTerpretability

### v5

@@ Abstract
Understanding the agent’s learning process, particularly the factors that contribute to its success or failure post-training, is crucial for comprehending the rationale behind the agent’s decision-making process. Prior methods clarify the learning process by creating a structural causal model (SCM) or visually representing the distribution of value functions. Nevertheless, these approaches have constraints as they exclusively function in 2D-environments or with uncomplicated transition dynamics. Understanding the agent’s learning process in complicated environments or tasks is more challenging. In this paper, we propose REVEAL-IT, a novel framework for explaining the learning process of an agent in complex environments. Initially, we visualize the policy structure and the agent’s learning process for various training tasks. By visualizing these findings, we can understand how much a particular training task or stage affects the agent’s performance in test. Then, a GNN-based explainer learns to highlight the most important section of the policy, providing a more clear and robust explanation of the agent’s learning process. The experiments demonstrate that explanations derived from this framework can effectively help in the optimization of the training tasks, resulting in improved learning efficiency and final performance.
@@ 1 Introduction
Based on the discussion about a good interpretability mechanism, we propose understanding the agent’s performance after training from a more fundamental perspective, i.e., the learning process of the policy and the sequences of training tasks. By focusing on the policy and tasks, we may effectively mitigate this problem. They possess a minimum of three evident benefits: (1) Both the training tasks and the policy provide well-organized and generalizable information, allowing it to work in diverse environments. (2) The information presented in the policy is aligned with the agent’s capacities. The limitation of the policy rests in our ability to extract intuitive explanations from the high-dimensional information it provides to understand the actions and behaviors of the agent. (3) When compared to SCM or counterfactual methods, which can not deal with the big and complicated problems that can not be seen in the real world, using structured training tasks as the goal can solve this problem. This can be further enhanced by extracting relevant information from the policy. At this location, the agent’s learning progress can serve as a temporal framework to enhance our comprehension of the agent’s behavior.
@@ 2 Related Works
Explainability for RL.  Puiutta & Veith (2020); Heuillet et al. (2020) categorize the explainable RL methods into two groups. The first group is “Post hoc”, which aims to provide explanations after the model execution, while the other group is “intrinsic approaches”, which inherently possess transparency. Many post hoc explanations, like the saliency map approach (Greydanus et al., 2017; Mott et al., 2019), are based on correlations. This means that conclusions drawn from correlations may not be accurate and do not answer questions about what caused what. You can use interpretable algorithms, like linear regression or decision trees, to develop intrinsic explanations. However, these algorithms may not be able to model complex behaviors (Puiutta & Veith, 2020) fully. Our method proposes a new mechanism in the problem of explainability for RL from the perspective of the agent’s learning process, i.e., it can help to understand how the agent is learning in a given task or environment.
@@
3 Preliminaries
Our framework comprises of two parts. One is the RL policy visualization module that shows the policy update in training and policy activation in the test. This module is based on fully connected neural networks’ well-established node-link diagram representation. The other is a GNN-based explainer, trained to analyze the visualized results in the first module and underly the critical update in the training phase. We can further explore the relationship between different training tasks and the contribution of a specific training task/phase to the final performance. Now, we provide more details on these two parts:
3.1 Reinforcement Learning and Visualization
@@
Reinforcement Learning. The problem of controlling an agent can be modeled as a Markov decision process (MDP), denoted by the tuple$\{\mathcal{S},\mathcal{A},P,R,\mu_{0},\gamma\}$, where $\mathcal{S}$ and $\mathcal{A}$ represent the set of state space and action space respectively, $P:\mathcal{S}\times\mathcal{A}\times S\rightarrow[0,1]$ is the transition probability function that yields the probability of transitioning into the next state $s_{t+1}$ after taking an action $a_{t}$ at state $s_{t}$, and $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function that assigns the immediate reward for taking an action $a_{t}$ at state $s_{t}$. $\mu_{0}:\mathcal{S}\rightarrow[0,1]$ is the probability distribution that specifies the generation of the initial state, and $\gamma\in[0,1]$ denotes the discount factor. We can use a deep RL algorithm to train the agent (e.g., DQN (Mnih et al., 2015), A2C (Mnih et al., 2016), SAC (Haarnoja et al., 2018b), PPO (Schulman et al., 2017), etc.). In the experiments of this paper, we use DQN and PPO as the fundational RL methods. Note that any online RL algorithm can be accepted.
@@
Visualization. The visualization should meet the following goals: summarize the unique properties of MLP networks in RL and reflect lessons learned from prior tasks. First, the visualization should handle networks large enough to solve complicated tasks. Second, the visualization should depict the network architecture with a node-link diagram. Third, the visualization should allow users to easily experiment with the input-output process of the network, allowing them to judge the network’s robustness to translational variance, rotational variance, and ambiguous input. Fourth, the visualization should allow users to view details on individual nodes, such as the activation level, the calculation being performed, the learned parameters (i.e., the node’s weights), and the numerical inputs and outputs. We choose to visualize the policy with a node-link diagram, and the only issue with a straightforward implementation of a node-link diagram is that large networks yield a dense mass of edges between layers. We base the visualization model on  Harley (2015), which can choose to display some or all of the weights connected to a certain node, and you can record and correspond to the changes in the value of a specific part of the weights during the training process.
@@ 3.2 Background on GNN-based Explainer
4 REVEAL-IT
In REVEAL-IT, assume that we first have a randomly drafted subtask sequence for the RL policy training. We formulate the training scheme of REVEAL-IT as in Alg. 1. We start from the conventional RL training process, i.e., the agent will explore and interact with the environment, trying to complete each sub-task. After the agent collects data, we can apply any online RL algorithm to update the control policy. Subsequently, we will evaluate the performance of the current policy every fixed number of environmental steps the agent taken. Iterate this procedure until the agent acquires a specific quantity of policy update data, at which point we can train the GNN explainer.
@@ 4 REVEAL-IT
1:Initialize: control policy $\pi_{\theta}$, RL replay buffer $\mathcal{B}\leftarrow\emptyset$, RL update data $\mathcal{D}\leftarrow\emptyset$, a set of training tasks $\mathcal{D}_{T}$, training tasks important score $y_{T}$;
@@
3:     if $t=1$ then $\triangleright$ Sample training task sequence
4:         Randomly sample training task sequence $Seq_{t}$ in $\mathcal{D}_{T}$;
5:     else
6:         Sample training tasks $Seq_{t}$ in terms of $y_{T}$;
7:     end if
8:     RL exploration and collect data to $\mathcal{B}$ following $Seq_{t}$; $\triangleright$ RL exploration
9:     Update $\pi_{\theta}$ using samples from $\mathcal{B}$;
10:     Visualize policy update and save update information into $\mathcal{D}$$\triangleright$ Visualization
11:     Train GNN to predict $y_{T}$ based on $\mathcal{D}$; $\triangleright$ Train GNN explainer
12:end for
The learning objective of GNN explainer on task. The overall goal of the GNN explainer is to learn to optimize the sequences of training tasks in terms of RL agent’s learning progress. To achieve this, the GNN explainer needs to be able to predict how much improvement an agent can acquire from a task, i.e., this GNN explainer needs to understand the visualized policy updates collected from previous iterations. Based on the discussion in Sec. 3.2, we denote the prediction task of our GNN explainer as $\Phi$: $G_{O}\rightarrow y$, where $y$ is the predicted improvement of the RL policy.
@@
Step 2. Highlight the important updates for explanations. As mentioned in Sec. 3.1, the control policy can be too complex to understand humans even with the visualized diagrams of the policy. Therefore, we also train the GNN explainer should highlight these important nodes and their related updates to humans to achieve a feasible explanation. The GNN explainer aims to partition $G_{O}$ into two subgraphs:
@@
where $G_{X}$ represents the explanatory subgraph reflecting the important edges (corresponding to policy updates), and $\Delta G$ is the remaining graph containing unimportant edges. When explaining a graph of visualized policy updates, we denote the optimal subgraph by $G_{X}^{m}$. When $G_{X}^{m}$ is used as input to $\Phi$, it can retain the original prediction. $G_{X}^{m}$ is optimal, as removing the features from it would result in a different prediction. We will apply GNN explainer to
@@ 5.2 Main Results
Figure 2: The first line depicts the sequential arrangement of sub-tasks that must be accomplished in a step-by-step manner to fully accomplish a given task. One of the sub-tasks is located on the left side of the second line. The first to third columns of the tree diagram illustrate the RL policy update process. The blue circles represent nodes in the neural network, while the connections between the circles represent weight updates. Thicker connections indicate larger updates in weight amplitude (selected by GNN explainer). The red circles in the tree diagram on the far right illustrate the specific policy nodes that are active during the evaluation process. The links here represent the revised weights throughout the training phase of this subtask. The orange squares indicate the portions of the policy that are common to several sub-tasks. We opt to depict the 8 interconnected nodes with the most significant weight adjustment, facilitating comprehension of the reinforcement learning policy’s learning process and highlighting the policy’s shared component more distinctly.
@@
The optimization of the training task sequences. We report the distribution of the sub-tasks in Fig. 4. As the training progresses, it becomes evident that the GNN-based explainer modifies the distribution of sub-tasks. Fig. 4 (left) shows the distribution of randomly sampled subtasks at the start of the training process. It is evident that the subtask “put” occurs most frequently. During the training, while examining the distribution of subtasks in Fig. 4 (middle left), task “put” remains the most common subtask, the frequencies of “look”, “pick”, and “find” have experienced a substantial increase. This demonstrates that these three qualities are important for the agent in ALFWorld. This aligns with our inherent comprehension of the environment, wherein the agent must initially acquire the ability to locate and retrieve the relevant item before it can proceed with future activities. As the training advances, the occurrence rate of these three task types declines proportionally (Fig. 4 (middle right)), suggesting that the agent already possesses the necessary skills. Training the agent with these tasks will not significantly enhance its performance. Fig. 4 (right) demonstrates that as the training advances, the training tasks increasingly emphasize the agent’s additional skills, which align with the original work’s final one or two stages. This indicates that the agent’s training is almost over, and the alterations in the distribution of the training task align with our comprehension of the agent’s learning process. It is evident that the explainer has comprehended and become proficient in the agent’s learning process, enabling them to effectively modify the training tasks.
@@
PPO+REVEAL-IT 1921.08 (0.90) 2104.88 (0.90) 1004.92 (0.90) -11.27 (0.90) 112.51 (0.90) 3148.64 (0.90)
@@
PG+REVEAL-IT 742.33 (0.90) 2253.70 (0.90) 975.04 (0.90) -13.21 (0.90) 70.66 (0.90) 1546.38 (0.90)

### v6

@@ Abstract
Understanding the agent’s learning process, particularly the factors that contribute to its success or failure post-training, is crucial for comprehending the rationale behind the agent’s decision-making process. Prior methods clarify the learning process by creating a structural causal model (SCM) or visually representing the distribution of value functions. Nevertheless, these approaches have constraints as they exclusively function in 2D-environments or with uncomplex transition dynamics. Understanding the agent’s learning process in complex environments or tasks is more challenging. In this paper, we propose REVEAL-IT, a novel framework for explaining the learning process of an agent in complex environments. Initially, we visualize the policy structure and the agent’s learning process for various training tasks. By visualizing these findings, we can understand how much a particular training task or stage affects the agent’s performance in the test. Then, a GNN-based explainer learns to highlight the most important section of the policy, providing a more clear and robust explanation of the agent’s learning process. The experiments demonstrate that explanations derived from this framework can effectively help optimize the training tasks, resulting in improved learning efficiency and final performance.
@@ 1 Introduction
Based on the discussion about a good interpretability mechanism, we propose understanding the agent’s performance after training from a more fundamental perspective, i.e., the learning process of the policy and the sequences of training tasks. By focusing on the policy and tasks, we may effectively mitigate this problem. They possess a minimum of three evident benefits: (1) Both the training tasks and the policy provide well-organized and generalizable information, allowing it to work in diverse environments. (2) The information presented in the policy is aligned with the agent’s capacities. The limitation of the policy rests in our ability to extract intuitive explanations from the high-dimensional information it provides to understand the actions and behaviors of the agent. (3) When compared to SCM or counterfactual methods, which can not deal with the big and complex problems that can not be seen in the real world, using structured training tasks as the goal can solve this problem. This can be further enhanced by extracting relevant information from the policy. At this location, the agent’s learning progress can serve as a temporal framework to enhance our comprehension of the agent’s behavior.
@@ 2 Related Works
Explainability for RL.  Puiutta & Veith (2020); Heuillet et al. (2020) categorize the explainable RL methods into two groups. The first group is “Post hoc”, which aims to provide explanations after the model execution, while the other group is “intrinsic approaches”, which inherently possess transparency. Many post hoc explanations, like the saliency map approach (Greydanus et al., 2017; Mott et al., 2019), are based on correlations. This means that conclusions drawn from correlations may not be accurate and do not answer questions about what caused what. However, these algorithms may not be able to model complex behaviors (Puiutta & Veith, 2020) fully. Our method proposes a new mechanism in the problem of explainability for RL from the perspective of the agent’s learning process, i.e., it can help to understand how the agent is learning in a given task or environment.
@@
3 Preliminaries on Reinforcement Learning and Visualization
@@
Our framework comprises of two parts. One is the RL policy visualization module that shows the policy update in training and policy activation in the test. This module is based on fully connected neural networks’ well-established node-link diagram representation. The other is a GNN-based explainer, trained to analyze the visualized results in the first module and underly the critical update in the training phase. We can further explore the relationship between different training tasks and the contribution of a specific training task/phase to the final performance. Now, we provide more details on these two parts: Reinforcement Learning. The problem of controlling an agent can be modeled as a Markov decision process (MDP), denoted by the tuple$\{\mathcal{S},\mathcal{A},P,R,\mu_{0},\gamma\}$, where $\mathcal{S}$ and $\mathcal{A}$ represent the set of state space and action space respectively, $P:\mathcal{S}\times\mathcal{A}\times S\rightarrow[0,1]$ is the transition probability function that yields the probability of transitioning into the next state $s_{t+1}$ after taking an action $a_{t}$ at state $s_{t}$, and $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function that assigns the immediate reward for taking an action $a_{t}$ at state $s_{t}$. $\mu_{0}:\mathcal{S}\rightarrow[0,1]$ is the probability distribution that specifies the generation of the initial state, and $\gamma\in[0,1]$ denotes the discount factor. We can use a deep RL algorithm to train the agent (e.g., DQN (Mnih et al., 2015), A2C (Mnih et al., 2016), SAC (Haarnoja et al., 2018b), PPO (Schulman et al., 2017), etc.). In this paper’s experiments, we use DQN and PPO as the fundational RL methods. Note that any online RL algorithm can be accepted.
@@
Strucutral visualization of the policy. The visualization should meet the following goals: summarize the unique properties of MLP networks in RL and reflect lessons learned from prior tasks. First, the visualization should handle networks large enough to solve complex tasks. Second, the visualization should depict the network architecture with a node-link diagram. Third, the visualization should allow users to easily experiment with the input-output process of the network, allowing them to judge the network’s robustness to translational variance, rotational variance, and ambiguous input. Fourth, the visualization should allow users to view details on individual nodes, such as the activation level, the calculation being performed, the learned parameters (i.e., the node’s weights), and the numerical inputs and outputs. We choose to visualize the policy with a node-link diagram, and the only issue with a straightforward implementation of a node-link diagram is that large networks yield a dense mass of edges between layers. We base the visualization model on  Harley (2015), which can choose to display some or all of the weights connected to a certain node, and you can record and correspond to the changes in the value of a specific part of the weights during the training process.
4 REVEAL-IT
In REVEAL-IT, assume that we first have a randomly drafted subtask sequence for the RL policy training. We start from the conventional RL training process, i.e., the agent will explore and interact with the environment, trying to complete each sub-task. After the agent collects data, we can apply any online RL algorithm to update the control policy. We can visualize the policy update with a node-link graph $G_{O}$ by comparing which part of the policy weights are updated. Subsequently, for every fixed number of environmental steps the agent takes, we can compute the true learning progress of the RL agent by evaluating the current policy. Then, we can construct the dataset for GNN explainer training based on the learning progress and graphs of policy updates $G_{O}$. We formulate the training scheme of REVEAL-IT as in Alg. 1 and introduce more details in REVEAL-IT as follows.
@@ 4.1 Background on GNN-based Explainer
1:Initialize: control policy $\pi_{0}$, RL replay buffer $\mathcal{B}\leftarrow\emptyset$, policy updated dataset $\mathcal{D}_{p}\leftarrow\emptyset$, a set of $N$ training tasks $\mathcal{D}_{task}$, randomly sampled sequence of training tasks $Seq_{0}$, the predicted learning progress $\{\mathcal{P}(task_{n},\pi_{0})=0|n=1,2,\cdots,N\}$;
@@
3:     $p$ = random();
4:     if $p<\epsilon$ then $\triangleright$ Sample training task sequence with $\epsilon-$greedy
5:         Randomly sample training task sequence $Seq_{t}$ from $\mathcal{D}_{task}$;
6:     else
7:         Sample training task sequence $Seq_{t}$ in terms of $\{\mathcal{P}(task_{n},\pi_{t})\}$;
8:     end if
9:     for $task_{i}$ in $Seq_{t}$ do
10:         RL agent explores and collects MDP data to $\mathcal{B}$ to complete $task_{i}$; $\triangleright$ RL exploration
11:         Use any online RL algorithm to update $\pi_{0}$ using samples from $\mathcal{B}$;
12:         Evaluate the current policy $\pi_{t}$ and compute $\mathcal{P}(task_{i},\pi_{t})$ following the Eq. 1;
13:         Visualize policy update information with graph $G_{O,t}$$\triangleright$ Visualization
14:         Save $(G_{O,t},\mathcal{P}(task_{i},\pi_{t}))$ into $\mathcal{D}_{p}$$\triangleright$ Construct dataset for GNN learning
15:     end for
16:     Train GNN predictor to minimize $|\hat{\mathcal{P}}-\mathcal{P}|^{2}$ based on $\mathcal{D}_{p}$;
17:     Train GNN explainer to partition $G_{O,t}$ into the optimal subgraph $G_{X,t}^{m}$;
18:end for
4.2 REVEAL-IT
The learning objective of GNN predictor. The overall goal of the GNN explainer is to learn to optimize the sequences of training tasks to improve the learning efficiency of the RL agent. We take inspiration from the learning progress in earlier work on curriculum learning (Narvekar et al., 2020; Ao et al., 2023), i.e., a higher learning progress signifies that the agent can acquire more knowledge regarding a specific training task, reflecting a higher learning efficiency. Therefore, in REVEAL-IT, we train the GNN predictor to learn to predict how much improvement the RL agent can have in a given task. Based on the discussion in Sec. 4.1, we denote the prediction task of our GNN explainer as $\Phi$: $G_{O}\rightarrow\hat{\mathcal{P}}$, where $\hat{\mathcal{P}}$ is the predicted improvement of the RL policy. Thus, in a given task $n$, we train the GNN predictor to minimize the error between $\hat{\mathcal{P}}$ and the true learning progress $\mathcal{P}$ defined as:
$\mathcal{P}(task_{n},\pi_{t})=\mathcal{R}(task_{n},\pi_{t})-\mathcal{R}(task_{n},\pi_{t-1}),$ (1)
where $\mathcal{R}(task_{n},\pi_{t})$ denotes the expected return that the RL agent achieves with policy $\pi_{t}$.
Train GNN explainer with RL agent simultaneously. In REVEAL-IT, we separately train a GNN explainer that helps to understand how an RL agent learns to complete tasks, i.e., the structural visualization of the policy allows us to observe which part of the policy weights is updated in one iteration, and the GNN explainer learns to highlight the most crucial updates related to the agent’s final success. One thing to note is that the training data used for the GNN explainer is based on the RL, which allows the GNN explainer to adapt to the learning process of the RL agent and can provide guidance during the training. More specifically, as the RL agent acquires MDP data from the interactions with the environment and learns from it, we can use any online RL algorithm to update the RL policy. After each policy update, we can obtain a structured update visualization of the policy by comparing the changes in the policy weight parameters and constructing the dataset for the GNN explainer based on the updated information. We can regard the visualized policy as a node-link graph, i.e., the updated weights can be viewed as the edges that connect with nodes in the graph. Thus, the learning objective of the GNN explainer can be described as “learn to find which part of the edges (updated weights) is more important for the success of the RL agent”. Now, we introduce more details of the GNN explainer to help the understanding.
@@ 4.2 REVEAL-IT
Step 2. Highlight the important updates for explanations. As mentioned in Sec. 3, the control policy can be too complex to understand humans, even with the visualized diagrams of the policy. Therefore, we also train the GNN explainer should highlight these important nodes and their related updates to humans to achieve a feasible explanation. The GNN explainer aims to partition $G_{O}$ into two subgraphs:
@@
where $G_{X}$ represents the explanatory subgraph reflecting the important edges (corresponding to policy updates), and $\Delta G$ is the remaining graph containing unimportant edges. When explaining a graph of visualized policy updates, we denote the optimal subgraph by $G_{X}^{m}$. When $G_{X}^{m}$ is used as input to $\Phi$, it can retain the original prediction. $G_{X}^{m}$ is optimal, as removing the features from it would result in a different prediction.
The GNN predictor and explainer serve distinct roles in REVEAL-IT. The former assesses the improvement of the training task on the RL policy performance, thereby increasing the learning efficiency, whereas the latter evaluates whether the GNN predictor comprehends the learning process of the RL agent by analyzing the correlation between “nodes linked to significant updates” and “the activated nodes during the test”. Simultaneously, by analyzing the graphs of the policy update by GNN explainer, we can ascertain the ability learned by the RL agent in a specific task and comprehend the significance of this capability in the RL agent’s success of the final task. REVEAL-IT is distinguished from other explainable RL methods by its independence from external data and structure, it learns to explain the learning process of the RL agent within the training task without imposing constraints on the environment or specific RL algorithms.
@@ 5.2 Main Results
Figure 2: Visualized important policy updates by GNN explainer. We provide a larger version in Appendix. D, and detailed analysis in section 5.3. In this figure, the first line depicts the sequential arrangement of sub-tasks that must be accomplished step-by-step to accomplish a given task fully. One of the sub-tasks is located on the left side of the second line. The first to third columns of the tree diagram illustrates the RL policy update process. The blue circles represent nodes in the neural network, while the connections between the circles represent weight updates. Thicker connections indicate larger updates in weight amplitude (selected by GNN explainer). The red circles in the tree diagram on the far right illustrate the specific policy nodes that are active during the evaluation process. The links here represent the revised weights throughout the training phase of this subtask. The orange squares indicate the portions of the policy that are common to several sub-tasks. We opt to depict the 8 interconnected nodes with the most significant weight adjustment, facilitating comprehension of the reinforcement learning policy’s learning process and highlighting the policy’s shared component more distinctly.
@@
The optimization of the training task sequences. We report the distribution of the sub-tasks in Fig. 4. As the training progresses, it becomes evident that the GNN-based explainer modifies the distribution of sub-tasks. Fig. 4 (left) shows the distribution of randomly sampled subtasks at the start of the training process. It is evident that the subtask “put” occurs most frequently. During the training, while examining the distribution of subtasks in Fig. 4 (middle left), task “put” remains the most common subtask, the frequencies of “look”, “pick”, and “find” have experienced a substantial increase. This demonstrates that these three qualities are important for the agent in ALFWorld. This aligns with our inherent comprehension of the environment, wherein the agent must initially acquire the ability to locate and retrieve the relevant item before it can proceed with future activities. As the training advances, the occurrence rate of these three task types declines proportionally (Fig. 4 (middle right)), suggesting that the agent already possesses the necessary skills. Training the agent with these tasks will not significantly enhance its performance. Fig. 4 (right) demonstrates that as the training advances, the training tasks increasingly emphasize the agent’s additional skills, which align with the original work’s final one or two stages. This indicates that the agent’s training is almost over, and the alterations in the distribution of the training task align with our comprehension of the agent’s learning process. It is evident that the explainer has comprehended and become proficient in the agent’s learning process, enabling them to modify the training tasks effectively.
@@
PPO+REVEAL-IT 1921.08 (0.90) $\uparrow$ 2104.88 (0.90) 1004.92 (0.90) $\uparrow$ -11.27 (0.90) 112.51 (0.90) $\uparrow$ 3148.64 (0.90) $\uparrow$
@@
PG+REVEAL-IT 742.33 (0.90) $\uparrow$ 2253.70 (0.90) 975.04 (0.90) -13.21 (0.90) $\uparrow$ 70.66 (0.90) $\uparrow$ 1546.38 (0.90) $\uparrow$
@@
Comparison with other interpretability frameworks In the experiments on ALFworld, we mainly focus on how REVEAL-IT can help in the RL training. However, it is necessary to evaluate the performance of our explanation mechanism, i.e., can other interpretability methods have better improvement for RL? Therefore, We add the comparison with other interpretability methods (GNNexplainer (Ying et al., 2019), MixupExplainer (Zhang et al., 2023)) and report the results in the Tab.ref 3. Due to the limited time, we can only compare with GNNexplainer and MixupExplainer in ALFworld, and we keep the hyperparameters the same as them in the released code.
Methods Avg. Pick Clean Heat Cool Look Pick2
REVEAL-IT (ours) 0.80 0.66 0.90 0.81 0.80 0.85 0.70
REVEAL-IT (with GNNexplainer) 0.64 0.47 0.70 0.61 0.69 0.70 0.55
REVEAL-IT (with MixupExplainer) 0.52 0.33 0.59 0.50 0.55 0.60 0.42
Table 3: The comparison with other interpretability methods in REVEAL-IT. In this ablation experiment, only the GNN explainer model was replaced, and other RL algorithms and parameters remained the same.